alones (alones):

oh this is a mindblown can somebody plz show me 219 asa a frraction cuz i just solved this equation/or what ever u call it...Given the function ƒ(x ) = 7x + 2, find x when ƒ(x ) = 31 the multiply choice is 1)4 2)33/7 3)29/7 this is *** for me even tho i know how to but right now i can't get any ideas to my head

OpenStudy (anonymous):

i just saw this yesterday lol

alones (alones):

ha,ha really well i don't know :)

OpenStudy (anonymous):

something to do with dividing the f(x)=31 by 7x but i forgot what you do with the 2 lol but you do something with the 2 before you divide i think lol ( i am no master )

alones (alones):

i feel stupid cuz i know how to solve it but i forgot maybe cuz i drank to much coffe :)

OpenStudy (anonymous):

lol yuck i hate coffee :)

OpenStudy (anonymous):

i also feel really dumb because i legit just had this question yesterday and now im like wth is this crap

alones (alones):

well yeah I solved this way f(x)=7931) +2 =217+2 =and here i got 219

alones (alones):

i need to make it asa fraction...

OpenStudy (anonymous):

7x+2= 9x 9x/31x maybeee but this seems to easy

OpenStudy (anonymous):

ohh as a fraction?

alones (alones):

ha,ha well i guess so how would u convert it to fraction

OpenStudy (anonymous):

girl please lol go buy some help from a qualified helper lol i hate fractions

OpenStudy (anonymous):

sorry lol

alones (alones):

that's fine i will :)

OpenStudy (anonymous):

could you maybe put 9x/31x because a fraction is just division???

OpenStudy (anonymous):

9x over 31x is what i meant

alones (alones):

eh u know what i think i'll just do long division :)

OpenStudy (anonymous):

lol good luck

OpenStudy (mathmate):

Hint: given f(x)=31, and f(x)=7x+2 Since they are both f(x), you can equate them: 31=7x+2 and solve for x.

alones (alones):

yeah that what i just did and i got again 219 :/ thx @mathmate ;)